# WebAccessVL: Making an Accessible Web via Violation-Conditioned VLM

### v1

@@ 3.2 Violation-conditioned VLM
Table 2: We show the evaluation metrics using open models and our violation-conditioned VLMs on the test set of WebAccessVL. Within the Baselines, the first block includes open APIs, the second contains open-source LLMs, and the third consists of open-source VLMs.
@@
Open API / Weights Baselines Raw Data - 5.335 - - -
GPT-4o 2.929 0.227 0.389 5.081
@@
$\hskip-123.76965pt\tilde{p}^{\text{NG}}_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{t}},{\bm{v}},{\bm{c}})=$
${\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\Bigr)}\Big/{\sum_{{\bm{u}}^{\prime}\in{\mathcal{V}}}\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}^{\prime},\gamma)\Bigr)}.$ (5)
4 Experiments
We benchmark various open API and open-weight models, followed by an evaluation of supervised fine-tuning using our proposed dataset. We then conducted a perceptual study to ensure the quantitative metrics aligned with human judgment. Finally, we defer the ablation study to the appendix Sec. A3.
Evaluation. A model is evaluated by its ability to improve accessibility while preserving the original design. We report four metrics to assess the accessibility and visual–structural quality:
  • Accessibility guidance violations $(\downarrow)$: Following IBM’s accessibility checker [19], we report the average number of detected WCAG2 violations per website.
  • Tree edit distance $(\downarrow)$: We compare the tree edit distance between the raw input HTML and the output HTML to validate HTML design consistency. Leveraging the Document Object Model (DOM) based HTML structure, we consider each HTML tag, such as <div>, <body>, <nav> as a tree node. We set the value of the cost of insert, update, and remove as one to penalize a structural disparity.
Baselines. To evaluate the accessibility quality of model-fixed websites, we selected 16 publicly open models, including both LLMs and VLMs:
  • LLMs: GPT-4o [31], GPT5 [32], Claude3.5 Sonnet [6], Gemini 1.5 Flash [38], Llama3.2 [11], Qwen2 [46], Qwen2.5 [47], and Qwen2.5 Coder [18].
  • VLMs: GPT4o Vision, DeepSeek VL2 [44], Janus Pro [8], Llama3.2 Vision [11], LLaVA1.5 [26], LLaVA1.6 [27], Qwen2 VL [40], and Qwen 2.5 VL [34].
Implementation details. Both LLM and VLM are adapted using LoRA [16]. Across all the models, we use Adam [20] to fine-tune the model for 100 epochs with a learning rate of $2\times 10^{-4}$. During sampling, we do top-1 sampling [13] and set the maximum sequence length to 2048. For negative guidance sampling, the guidance scale is set to 1.1 for LLaVA 1.6 and 1.05 for Llama 3.2 Vision. Input images are resized to $640\times 360$. All experiments fit in a single 48GB Nvidia A6000 GPU. Additional implementation details are provided in Appendix A4.
Figure 3: The percentage improvement in violation for open API/weights models compared to the raw HTML of the test set.
4.1 Comparison with Open API / weights models
To understand the difficulty of this task, we evaluate the performance of publicly available API-based and open-weight models using prompting. For fairness, all models are tested with the same instruction prompt as used in our method (see Appendix A4 for details). In Tab. 2, we report the results across LLMs and VLMs. We also present our fine-tuned models with negative guidance, initialized from LLaVA 1.6 [27] and Llama 3.2 Vision [11].
Results on LLMs. Claude3.5 Sonnet and GPT-4o are the best performing open APIs. They achieve moderate gains in improving accessibility with some degree of content/design preservation. While GPT-5 achieves the lowest number of violations, it significantly alters the website design, as indicated by its low structure accuracy and high tree edit distance. For Gemini 1.5 Flash [38], it slightly increased the number of violations. When comparing to open-source LLMs, such as Llama 3.2 or Qwen, we find that they, in general, perform worse than the open-API models.
Results on VLMs. For models that take in visual inputs, we further provide the rendering of the HTML. For GPT-4o Vision, it achieved a lower number of violations compared to its pure LLM counterpart and achieved an improved structure accuracy. Other models are able to fix some violations, except DeepSeek VL2 small [44] that slightly increases the violation count. We also visualized the percentage improvement in violation from the raw data in Fig. 3.
Results on Ours. As shown in Tab. 2, our method convincingly outperforms the baseline models. Our approach significantly lowers the number of violations to 0.5 violations per website on average. Our method also produces more accurate alt text and better preserves the website structure. Our method on Llama 3.2 Vision and LLaVA 1.6 shows an improvement in the number of violations of 91.8% and 90.5%. For structural accuracy, our models outperform all open-source models. The slightly lower SSIM compared to GPT-4o is due to adjustments of text contrast, which may reduce the SSIM and tree edit distance but improve accessibility significantly. Overall, these results demonstrate the importance of the proposed WebAccessVL dataset with SFT for developing automated systems to improve accessibility. We now study SFT in more detail.
Figure 4: The percentage improvement in violation for SFT models compared to the raw HTML of the test set.
Figure 5: The percentage of vision/language violations out of all the violations from each LLM and VLM.
Table 3: Results of SFT the violation-conditioned VLM with different pre-trained models on WebAccessVL. The results of LLMs are in the first block, and those of VLMs are in the second block.
@@ 4.1 Comparison with Open API / weights models
SFT Models Llama 3.2 1.119 0.208 0.255 3.104
Qwen 2 7B 0.896 0.217 0.355 2.776
Qwen 2.5 7B 0.743 0.220 0.310 2.926
Qwen 2.5 Coder 7B 0.579 0.221 0.304 2.947
LLaVA 1.5 0.810 0.244 0.313 2.989
LLaVA 1.6 0.527 0.262 0.307 3.049
Qwen2 VL 0.598 0.220 0.317 2.938
Qwen2.5 VL 0.700 0.261 0.330 2.825
Llama 3.2 Vision 0.451 0.260 0.331 3.004
@@
4.2 Comparison with SFT models
We examine the effect of supervised fine-tuning (SFT) on our violation-conditioned model with different pre-trained models. Results are shown in Tab. 3, where we observe that all models fine-tuned on WebAccessVL show a reduction in the number of violations.
Importance of visual information in SFT. In Fig. 4, we plot the percentage of improvement of violation # relative to the raw data. We observe that on average, VLMs tend to improve more over LLMs. For instance, the Llama 3.2 model with and without vision has a large gap in the percentage of improvement. From Tab. 3, we see that Llama 3.2 (LLM) has an average of 1.119 violations per HTML code, which drops to 0.451 when visual input is added. Furthermore, the caption-image CLIP score increases from 0.208 to 0.260, highlighting the benefit of visual information. In other words, the VLM variant better preserves the rendered structure, as indicated by higher structural accuracy and a lower tree edit distance. A similar trend is also observed between Qwen 2.5 7B and its VLM variant.
Analysis on vision vs. language violations. To further investigate whether violation-conditioned VLMs outperform LLMs, we plot the proportion of vision and language violations relative to the total violations made by each model in Fig. 5. Across the same backbone, we observe that Qwen 2.5 VL exhibits over 25% fewer vision-related violations than Qwen 2.5. This decrease is primarily due to improved handling of text contrast and image captioning in alt attributes. On average, 54.5% of the violations for LLMs are related to vision tasks, compared to 46.3% for VLMs. This reduction of 8.1% demonstrates that VLMs with visual input are more effective at addressing vision violations than pure LLMs. Critically, LLMs often fail to detect insufficient text contrast, as they rely solely on textual HTML content and rarely pick up on visual styles such as foreground-background color pairs. We hypothesize that as VLMs can directly inspect rendered screenshots, this enables them to correct vision violations more reliably.
Ablation of negative guidance. In Tab. 3, we report results with and without negative guidance. We observe that incorporating negative guidance into our violation-conditioned VLMs further reduces the number of violations compared to their counterparts without negative guidance. Specifically, the violation count for LLaVA 1.6 decreases from 0.527 to 0.509 with standard deviation of 0.0012, and for Llama 3.2 Vision, it drops from 0.451 to 0.439 with standard deviation of 0.0017, both averaged over three random seeds. At the same time, the other metrics remain comparable to those of the version without negative guidance.
@@ 4.2 Comparison with SFT models
Figure 6: Qualitative results with Ours, Claude 3.5 Sonnet [6], and ChatGPT4o Vision [31]. Yellow arrows highlight the fixed components, and red arrows show inaccessible parts.
4.3 Qualitative Results
In Fig. 6, we present qualitative comparisons on in-the-wild websites to show our model’s ability to generalize beyond the WebAccessVLdataset. In the first row, we observe that our model fixed the header, main background image, button, and footer contrast and put a gray bar at the bottom. However, Claude 3.5 did not change the header contrast but did change the gray bar at the bottom to enhance readability. GPT-4o generated a text close to the button in the main image, which makes it hard to read, and removed one of the products.
In the second example, our model fixed the contrast of the header, top three icons, green buttons, and the footer to improve readability. In contrast, Claude 3.5 did not fix the contrast of the header. Also, the footer color has insufficient contrast. GPT-4o missed two products and failed to put the footer in the right location. Overall, our model can fix contrasts of visual assets to improve visibility and make minimal changes to the website structure/style.
4.4 Perceptual Study
Study design. We conducted a perceptual study to assess how well the modified HTML files preserve both the website’s content and visual style. We compare against Claude 3.5 and GPT-4o, which achieved a relatively low number of accessibility violations while maintaining overall design fidelity, as observed in Tab. 2. Although GPT-5 achieves a low violation count, it makes large changes to the website’s design.
Each participant is shown the reference webpage along with outputs from our method and two baselines, Claude 3.5 and GPT-4. Specifically, participants were asked, “Which website (A, B, or C) best maintains the content and visual style of the reference image? In the event of a tie, please select the one with the better visual design.” Participants may select only one option for each question or leave the question empty. We gathered statistics on 13 questions and received responses from 31 participants.
Statistical hypothesis setup. We analyze user preference data using a chi-squared test of independence to determine whether the distribution of selections differs significantly across the three methods. Let the three methods be denoted as $A$, $B$, and $C$ (AA is Ours, $B$ is Claude 3.5, and $C$ is GPT-4o). We first perform a chi-squared test of independence on the following hypothesis (More details in Sec. A5):
@@ 4.4 Perceptual Study
Results and statistical analysis. We aggregate selections over all 13 questions with 401 valid responses, with counts:
$(n_{A},n_{B},n_{C})=(248,\;54,\;99).$ (9)
The chi-squared test of independence on the 3×1 contingency table yielded
$\chi^{2}_{\mathrm{global}}(2,\,N=401)=154.27,\quad p<0.001.$ (10)
@@
$\chi^{2}_{A>B}(1,\,N_{A,B}=302)=130.01,\quad p_{A>B}<10^{-16},$ (11)
$\chi^{2}_{A>C}(1,\,N_{A,C}=347)=64.00,\quad p_{A>C}<10^{-15}.$ (12)
Both $p_{A>B}$ and $p_{A>C}$ are below $\alpha_{\mathrm{adj}}=0.0167$. These pairwise results demonstrate that participants significantly selected our method over both Claude 3.5 and GPT-4o, i.e., our method better preserves the website’s content and visual style. This conclusion is in agreement with the quantitative structural metrics (Struct. Acc and Tree Edit Dist.) presented in Tab. 2.
@@ 5 Conclusion
In this work, we formulate the task of improving a website’s accessibility as an image-conditioned program synthesis problem. We construct a dataset consisting of paired HTML files that have been manually modified to comply with WCAG2. We propose training VLMs in a violation-conditioned manner to address this task. Empirical results show that our method reduces the average number of violations from 5.34 to 0.44 per website on the test set. Compared to GPT-5, our approach achieves a 73.9% reduction in violations. We also show that LLMs are suboptimal in understanding visual content when relying solely on HTML code. This work is a first step toward automating web accessibility. To support further research, we will release both the code and the dataset.
@@
  • We conduct more ablation studies on the training module and training set size in Sec. A3.
@@
We show additional qualitative results using our violation-conditioned model, Claude 3.5 Sonnet [6], and ChatGPT4o Vision [31] in Fig. A1 to Fig. A6.
@@
Figure A1: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours modifies the contrast between the header, a button, and the descriptions in each card, while Claude 3.5 fails to put valid image links. Moreover, ChatGPT4o could not correct the contrast of the header. Please zoom in for details.
@@
Figure A2: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours corrects the text contrast in the navigation bar, descriptions in each image, and side bars. Claude 3.5 does not have appropriate contrast on the title of each card and is missing the thumbnails. ChatGPT4o also has not completely fixed the text contrast in the left side bar, and generates a different website layout ratio. Please zoom in for details.
@@
Figure A3: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours successfully put an accessible contrast in buttons and the texts. Claude 3.5 and chatGPT4o failed to put an appropriate color contrast on the buttons. Please zoom in for details.
@@
Figure A4: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows that it generates an accessible contrast on the header, description, and footer. Moreover, ours preserves the website style by putting all four images in a row, while Claude 3.5 and chatGPT4o could not preserve the original design. Please zoom in for details.
@@
Figure A5: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it can keep the style while correcting inaccessible elements by changing their color contrast values. Claude 3.5 failed to preserve the original web design by making thicker header and footer thicker. Also, it could not put the correct image URLs. Please zoom in for details.
@@
Figure A6: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it fixes text contrast in the header, image cards, and the footer by following the original web style. Claude 3.5 did not preserve the original style by making larger selection boxes and missing images. ChatGPT4o failed to fix the proper text and button contrasts to comply with the accessibility guideline. Please zoom in for details.
@@
Qwen2 VL QV 1.123 (79.0) 0.202 0.318 2.893
QKVO 1.127 (78.9) 0.243 0.309 3.101
QKVO+ FC + LN 0.621 (88.4) 0.255 0.325 2.794
QKVO + VP + VR 0.598 (88.8) 0.220 0.317 2.938
Qwen2.5 VL QV 3.491 (34.6) 0.164 0.632 1.554
QVKO 1.625 (69.6) 0.220 0.382 3.210
QKVO + FC + LN 1.672 (68.7) 0.229 0.359 2.198
QKVO + VP + VR 0.700 (86.9) 0.261 0.330 2.825
Llama 3.2 Vision QV 0.548 (89.7) 0.249 0.295 2.995
QKVO 0.582 (89.1) 0.242 0.302 2.996
QKVO + MLP 0.885 (83.4) 0.260 0.356 2.884
QV + DP + UP + GateP 0.451 (91.6) 0.260 0.331 3.004
LLaVA 1.5 QV 0.958 (82.0) 0.234 0.326 2.940
QVKO 1.349 (74.7) 0.247 0.395 3.076
QKVO + Vision proj 0.810 (84.8) 0.244 0.313 2.989
QKVO+ FC + LN 1.182 (77.8) 0.232 0.450 2.488
LLaVA 1.6 QV 0.627 (88.2) 0.248 0.310 2.989
QVKO 0.510 (90.4) 0.259 0.293 3.097
QKVO + Vision proj 0.618 (88.4) 0.246 0.311 3.032
QKVO+ FC + LN 0.527 (90.1) 0.262 0.307 3.049
Table A1: Results of ablating different training modules from the test set of WebAccessVL.
Ablating trainable modules for SFT VLMs. We conduct ablation studies to analyze the impact of different training modules on our model’s performance. In Tab. A1, we present results across multiple open-source VLMs and their training module variations. The results indicate that Query, Key, Value, and Out Projection generally enhance Caption-Image CLIP similarity but may increase violations. Incorporating fully connected layers, layer norms, or structured projections helps reduce violations while maintaining alignment. Among the variations, we report the one with the fewest violations in Tab. 3.
Ablating training sample size for SFT VLMs. We investigate the effect of training sample size in Tab. A2. As the number of training samples increases, violations generally decrease, though not strictly, suggesting that models can achieve better alignment with accessibility guidelines when trained with more appropriate data. For Caption-Image CLIP similarity, we observe an increasing trend in Llama 3.2 Vision [11] and LLaVA 1.5 [26], while other models do not exhibit a consistent trend.
[…]
Qwen2 VL 100 1.364 (74.4) 0.217 0.287 2.835
200 0.903 (83.1) 0.221 0.304 2.925
300 0.836 (84.3) 0.209 0.289 3.034
400 0.888 (83.4) 0.208 0.316 2.977
Qwen2.5 VL 100 2.153 (59.6) 0.212 0.414 2.182
200 3.054 (42.8) 0.162 0.612 1.517
300 3.276 (38.6) 0.137 0.666 1.224
400 2.102 (60.6) 0.188 0.479 2.337
Llama 3.2 Vision 100 0.674 (87.4) 0.227 0.237 2.989
200 0.602 (88.7) 0.243 0.320 2.847
300 0.574 (89.2) 0.246 0.328 2.967
400 0.589 (89.0) 0.260 0.318 3.010
LLaVA 1.5 100 1.065 (80.0) 0.234 0.253 3.786
200 1.060 (80.1) 0.239 0.286 3.618
300 1.278 (76.0) 0.231 0.362 3.168
400 0.870 (83.7) 0.242 0.315 3.025
LLaVA 1.6 100 0.895 (83.2) 0.211 0.308 3.135
200 0.950 (82.2) 0.151 0.290 3.101
300 0.698 (86.9) 0.226 0.320 3.047
400 0.597 (88.8) 0.202 0.317 2.977
Table A2: Results of ablating different training sizes from the test set of WebAccessVL.
A4 Implementation Details & Prompt Ablation
@@
Training details. The trainable module and the trainable parameter percentage of each model are provided in Tab. A3. Across all models, we train LoRA [16] with a rank of $r=8$ and $\alpha=16$, resulting in a scaling factor of $\alpha/r=2$. The dropout rate is set to 0.05. Due to memory constraints, we use gradient accumulation to enable training with a larger batch size. Specifically, we train Llama 3.2 Vision [11] with a batch size of 32 and LLaVA 1.6 [27] with a batch size of 24, both using negative guidance. The number of epochs for all the training runs is 100. During training, the input text prompt is:

### v2

@@ 4.2 Main Results
Table 1: We show the evaluation metrics using the best open models and our violation-conditioned VLMs on the test set of WebAccessVL. We present open APIs, open-source LLMs, and open-source VLMs.
@@
Raw Data - 5.335 - - -
Open API GPT-4o 2.929 0.227 0.389 5.081
@@
Impact on disability groups. We further analyze how our method impacts specific disability groups following the categorization by Droutsas et al. [12] shown in Tab. 2. Overall, our method fixes 98.2% of violations affecting
Table 2: Violation fixes by disability group [12] on 1,000 test samples.
Group Before After Fix Rate
Blind/Low-vision 2,467 45 98.2%
Motor/Keyboard 1,334 24 98.2%
Low vision/Dyslexia 1,525 142 90.7%
Overall 5,326 211 96.0%
blind/low-vision users (2,467 violations: missing alt text, ARIA labels, page titles) and motor-impaired users (1,334 violations: keyboard navigation, skip links, landmark structure). Text contrast violations affecting dyslexia and low-vision users have a 90.7% fix rate (1,383/1,525). This is lower because some fixes require changing the original design. The remaining 4% involve rare custom ARIA roles that appear in <1% of training data.
4.3 VLMs vs. LLMs Comparison
We compare LLMs and VLMs using the same base architecture to evaluate the importance of visual information. Tab. 3 shows results for GPT-4o, Llama 3.2, and Qwen 3 in text-only (LLM) and vision-enabled (VLM) configurations, with and without SFT.
Table 3: Comparison between text-only LLMs and VLMs with and without supervised fine-tuning on WebAccessVL.
Base Model VLM SFT # Violation $\downarrow$ Caption-Img CLIP $\uparrow$ Struct. Acc. $\uparrow$ Tree Edit Dist. $\downarrow$
Raw Data - - 5.335 - - -
GPT 4o 2.929 0.227 0.389 5.081
2.918 0.250 0.408 4.939
Llama 3.2 4.382 0.173 0.020 14.097
4.765 0.165 0.019 15.474
1.119 0.208 0.255 3.104
0.451 0.260 0.331 3.004
Qwen 3 5.185 0.216 0.021 0.567
5.227 0.235 0.022 0.412
2.670 0.176 0.021 6.661
0.604 0.239 0.019 11.807
VLMs and LLMs are comparable without fine-tuning. For GPT-4o, adding vision capabilities gives marginal improvements in the zero-shot setting (2.929 $\rightarrow$ 2.918 violations), with slight gains in structural metrics. Generic vision-language pre-training does not equip models to visually diagnose accessibility issues. While the model can “see” the page, it does not know which visual patterns indicate violations.
SFT improves the importance of visual input. For Llama 3.2, SFT reduces violations from 4.382 to 1.119 for text only, but adding vision reduces this to 0.451, a 60% additional improvement. The VLM version achieves better structural preservation (0.331 vs 0.255), showing that visual understanding helps make more targeted modifications. That is, fine-tuning teaches the model to extract accessibility relevant visual features.
Visual input helps vision violations. Qwen3 shows a clear pattern; without SFT (zero-shot) performance is nearly identical between LLM and VLM (5.185 vs. 5.227). With SFT, the Qwen3 achieves 77% fewer violations (2.670 vs. 0.604). VLMs excel at vision-dependent issues (contrast, alt text quality) while performing comparably on structural issues (ARIA landmarks, semantic HTML). Lastly, both LLM and VLM show similarly low structural accuracy post SFT (0.021 and 0.019, respectively), indicating that the accessibility modifications required exceed the strict SSIM threshold regardless of whether visual understanding is used.
4.4 Ablation on Violation-Awareness
We conduct ablations to evaluate the contribution of our violation conditioning approach. Tab. 4 compares four configurations: (a) pre-trained model with zero-shot prompting, (b) standard supervised fine-tuning (SFT), (c) SFT with violation conditioning (VC), and (d) VC with iterative checker in the loop refinement.
Table 4: Comparison of pre-trained versus SFT on WebAccessVL, with and without violation-conditioned prompting (VC) and checker-in-the-loop refinement (Loop).
Base Model Method # Violation $\downarrow$ Caption-Img CLIP $\uparrow$ Struct. Acc. $\uparrow$ Tree Edit Dist. $\downarrow$
Raw Data - 5.335 - - -
Llama 3.2 Pre-train 4.765 0.165 0.019 15.474
SFT 0.451 0.260 0.331 3.004
+ VC 0.337 0.241 0.018 11.687
@@ 4.4 Ablation on Violation-Awareness
Gemma 3 Pre-train 4.427 0.233 0.021 0.313
SFT 0.440 0.238 0.018 12.434
+ VC 0.352 0.239 0.019 11.747
[…]
Violation conditioning is effective. Adding VC reduces violations by 25–30% across models (Llama: 0.451 $\rightarrow$ 0.337, LLaVA: 0.706 $\rightarrow$ 0.517). For Llama 3.2, structural accuracy decreases from 0.331 to 0.018, while for LLaVA 1.6 it drops slightly from 0.020 to 0.019. This suggests that violation conditioning leads to more comprehensive accessibility fixes that exceed the strict SSIM threshold of 0.9, though the qualitative results and perceptual study confirm these changes preserve visual appearance. With conditioning, models make targeted edits focused on elements flagged in the violation report.
Checker-in-the-loop fixes induced violations. Iterative refinement gives a small but consistent gain (Llama: 0.337 $\rightarrow$ 0.244, LLaVA: 0.517 $\rightarrow$ 0.366, Gemma: 0.352 $\rightarrow$ 0.211). Structural accuracy remains stable, showing that the iterative process focuses on fixing violations without introducing additional structural changes. Note, single-pass models may introduce new violations while fixing existing ones. For example, when attempting to fix contrast issues by modifying the background and text colors, the model may simultaneously update both the background and text to be dark, i.e., creating a new low-contrast violation. The checker-in-the-loop catches these violations (induced by the first pass) in the second round. Most samples converge in 2 to 3 iterations. Finally, we observe consistent improvement across several models, demonstrating the generality of our violation conditioning approach.
@@
Figure 4: Qualitative results with Ours, Claude 3.5 Sonnet [6], and GPT-5 [33]. Yellow arrows highlight the fixed components, and red arrows show inaccessible parts.
4.5 Qualitative Results & Perceptual Study
Qualitative results. Fig. 4 presents qualitative comparisons of in-the-wild websites to show our model’s ability to generalize beyond the WebAccessVL dataset. In the first example, our model corrects the header text contrast, adjusts the background image overlay, fixes button colors, and improves footer legibility, all while preserving the original layout. Claude 3.5 only partially addresses the contrast issues: the hero text and primary navigation links remain inaccessible. GPT-5 attempts a more aggressive fix but misaligns the header over the background image and introduces new low-contrast button colors.
In the second example, our model fixes contrast violations in the header, icon labels, call-to-action buttons, and footer while keeping the page structure intact. Claude 3.5 leaves the header and footer with insufficient contrast. GPT-5 restructures the page significantly, dropping destination cards and repositioning the footer, consistent with its low structural accuracy in Tab. 1. Overall, our model fixes contrast violations while preserving the visual layout and content of the original website.
Perceptual study design. We conducted a perceptual study to assess how well the modified HTML files preserve both the website’s content and visual style. We compare against Claude 3.5 and GPT-5. Each participant is shown the reference webpage along with outputs from our method and two baselines, Claude 3.5 and GPT-5. Participants were asked, “Which website (A, B, or C) best maintains the content and visual style of the reference image? If tied, select the one with the better visual design.” Participants may select only one option for each question or leave the question empty. We collected 13 questions with responses from 30 participants.
Statistical hypothesis setup. We analyze user preference data using a chi-squared test of independence to determine whether the distribution of selections differs significantly across the three methods. Let the three methods be denoted as $A$, $B$, and $C$ (AA is Ours, $B$ is Claude 3.5, and $C$ is GPT-5). We first perform a chi-squared test of independence on the following hypothesis (More in Sec. A6):
@@ 4.5 Qualitative Results & Perceptual Study
Results and statistical analysis. We aggregate selections over all 13 questions with 390 valid responses, with counts:
$(n_{A},n_{B},n_{C})=(195,\;76,\;119).$ (10)
The chi-squared test of independence on the $3\times 1$ contingency table yielded
$\chi^{2}_{\mathrm{global}}(2,\,N=390)=55.86,\quad p<0.001.$ (11)
@@
$\chi^{2}_{A>B}(1,\,N_{A,B}=271)=26.13,\quad p_{A>B}<10^{-6},$ (12)
$\chi^{2}_{A>C}(1,\,N_{A,C}=314)=9.20,\quad p_{A>C}<0.002.$ (13)
Both $p_{A>B}$ and $p_{A>C}$ are below $\alpha_{\mathrm{adj}}=0.0167$. Participants prefer our method over both Claude 3.5 and GPT-5. Indicating that our method better preserves the website’s content and visual style.
@@ 5 Conclusion
In this work, we formulate the task of improving website accessibility as an image-conditioned program synthesis problem. We construct WebAccessVL, a dataset of 1,500 paired HTML files manually modified to comply with WCAG2. We propose training VLMs in a violation-conditioned manner, where the model conditions on parsed violation reports describing specific accessibility issues detected in the input HTML. This conditioning enables the model to focus on fixing the violations in the report. Empirical results show that our method reduces violations from 5.34 to 0.211 per website with Gemma 3 and 0.244 with Llama 3.2 Vision, a 96.0% reduction from raw data. Compared to GPT-5 (1.68 violations), our approach achieves an 87% reduction while maintaining design fidelity, with 90% structural accuracy vs. GPT-5’s 0.5%. Through systematic comparisons, we demonstrate that VLMs significantly outperform text-only LLMs, with fine-tuned VLMs achieving 60–77% fewer violations than their LLM counterparts by leveraging visual understanding of rendered pages. This work represents an important step toward automating web accessibility. To support further research, we will release both the code and the dataset.
@@
  • We conduct ablation studies on violation conditioning strategies and negative guidance in Sec. A3.
  • We analyze the computational efficiency of our checker-in-the-loop approach in Sec. A4.
@@
We show additional qualitative results using our violation-conditioned model, Claude 3.5 Sonnet [6], and GPT-5 [33] in Fig. A1–Fig. A2 and Fig. A3–Fig. A6.
@@
Figure A1: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours modifies the contrast between the header, a button, and the descriptions in each card, while Claude 3.5 fails to put valid image links. Moreover, GPT-5 could not correct the contrast of the header. Please zoom in for details.
@@
Figure A2: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours corrects the text contrast in the navigation bar, descriptions in each image, and sidebars. Claude 3.5 does not have appropriate contrast on the title of each card and is missing the thumbnails. GPT-5 also has not completely fixed the text contrast in the left sidebar. Please zoom in for details.
@@
Figure A3: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours successfully put an accessible contrast in buttons and the texts. Claude 3.5 and GPT-5 failed to put an appropriate color contrast on the buttons. Please zoom in for details.
@@
Figure A4: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows that it generates an accessible contrast on the header, description, and footer. GPT-5 fails to put a consistent blue color in the button. Moreover, ours preserves the website style by putting all four images in a row, while Claude 3.5 could not preserve the original design. Please zoom in for details.
@@
Figure A5: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it can keep the style while correcting inaccessible elements by changing their color contrast values. GPT-5 failed to put a correct contrast to buttons. Claude 3.5 failed to preserve the original web design by making the header and footer thicker. Also, it could not put the correct image URLs. Please zoom in for details.
@@
Figure A6: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it fixes text contrast in the header, image cards, and the footer by following the original web style. Claude 3.5 did not preserve the original style by making larger selection boxes and missing images. GPT-5 failed to put a correct contrast on the button and the footer. Please zoom in for details.
@@
Violation conditioning: binary vs. report. We compare two conditioning strategies for guiding the VLM: (1) Binary, which conditions the model on a binary indicator of whether violations exist (zero/non-zero), and (2) Report, which conditions on the full structured violation report $\bm{c}$} containing detailed violation types, messages, and locations. As shown in Tab. A1, conditioning on the full violation report (Report) achieves fewer violations than binary conditioning (LLaVA 1.6: 0.366 vs. 0.509; Llama 3.2 Vision: 0.244 vs. 0.439), but results in substantially more aggressive structural modifications, as reflected by much higher tree edit distance (LLaVA 1.6: 11.141 vs. 3.053; Llama 3.2 Vision: 11.642 vs. 3.021). Instead, binary conditioning makes more conservative edits with lower tree edit distance, limiting unintended structural side-effects. Therefore, we use binary violation conditioning in our main experiments to balance accessibility improvement with structural fidelity.
@@
LLaVA 1.6 Binary 0.509 0.242 0.300 3.053
Report 0.366 0.236 0.910 11.141
Llama 3.2 VL Binary 0.439 0.260 0.331 3.021
Report 0.244 0.241 0.900 11.642
Table A1: Ablation study on violation conditioning strategies. We compare conditioning on violation count (zero/non-zero) versus full violation report (detailed violation types and locations). Violation count conditioning achieves better design preservation (higher SSIM, lower tree edit distance) while maintaining low violation counts, demonstrating a favorable balance between accessibility improvement and structural fidelity.
Negative guidance ablation. We analyze the impact of negative guidance sampling combined with the checker-in-the-loop mechanism across three VLMs. As shown in Tab. A2, negative guidance consistently reduces violations across all models. For LLaVA 1.6, violations decrease from 0.557 (Pass 1 without negative guidance) to 0.517 (Pass 1 with negative guidance), and further to 0.366 after the checker-in-the-loop refinement. Similar trends are observed for Llama 3.2 Vision (0.364 $\rightarrow$ 0.337 $\rightarrow$ 0.244) and Gemma 3 (0.370 $\rightarrow$ 0.352 $\rightarrow$ 0.234). Notably, the checker-in-the-loop mechanism provides additional refinement with minimal overhead, as only 10.7%–18.9% of files require Pass 2 (see Tab. A3). The combination of negative guidance and iterative refinement achieves the best performance, reducing violations by 33.0%–36.8% compared to the initial Pass 1 without negative guidance, while maintaining high structural accuracy across all models.
Model W/O Negative Guidance W/ Negative Guidance
W/O Loop W/ Loop W/O Loop W/ Loop
LLaVA 1.6 0.557 0.381 0.517 0.366
Llama 3.2 Vision 0.364 0.285 0.337 0.244
Gemma 3 0.370 0.259 0.352 0.234
Table A2: Ablation study on negative guidance sampling with checker-in-the-loop. We compare violation counts with and without negative guidance across multiple passes. W/O Loop shows results after the first generation, while W/ Loop shows the best result after up to $K=2$ passes. Negative guidance consistently reduces violations, particularly in the first pass, demonstrating its effectiveness in improving accessibility compliance.
A4 Computational Efficiency of Checker-in-the-Loop
Model Files Requiring Second Pass
LLaVA 1.6 18.9%
Llama 3.2 Vision 11.4%
Gemma 3 10.7%
Table A3: Computational overhead of checker-in-the-loop. We report the percentage of files requiring a second pass in our iterative refinement process (up to $K=2$ passes).
To demonstrate that our checker-in-the-loop approach achieves substantial improvement with minimal computational overhead, we analyze the percentage of files requiring a second pass across different models. As shown in Tab. A3, only 10.7%–18.9% of files require Pass 2 across all models, with Gemma 3 requiring the fewest at 10.7%. This indicates that the vast majority of websites (over 81%) converge to acceptable accessibility standards after just the first generation. Combined with the violation reduction shown in Tab. A2, these results demonstrate that our iterative refinement mechanism provides significant quality improvements while imposing minimal additional computational cost in practice.
A5 Implementation Details & Prompt Ablation
@@
Qwen 3 VL QKVO, Vision Proj, Vision Resampler 6.08
@@
Training details. The trainable module and the trainable parameter percentage of each model are provided in Tab. A4. Across all models, we train LoRA [19] with a rank of $r=8$ and $\alpha=16$, resulting in a scaling factor of $\alpha/r=2$. The dropout rate is set to 0.05. Due to memory constraints, we use gradient accumulation to enable training with a larger batch size. Specifically, we train Llama 3.2 Vision [14] with a batch size of 16, LLaVA 1.6 [27] with a batch size of 24, and Gemma 3 with a batch size of 13, all using negative guidance. The number of epochs for all the training runs is 100. During training, the input text prompt is: